OpenStudy (anonymous):

find vertex of y=3(x-3)^2+4

OpenStudy (amistre64):

since it is already in vertex form ... y = a (x-Vx) + Vy where the vertex is (Vx, Vy)

OpenStudy (anonymous):

so it would be (-3,-4) x=3

OpenStudy (raden):

i think it should has power 2 for in bracket, right @amistre64 ?

OpenStudy (anonymous):

\[y=3(x-3)^{2}+4\]

OpenStudy (anonymous):

y=a(x+b) + c your turning point will simply be (-b,c) the b value in your example is -3 so it will be positive 3, just remember it's the opposite.

OpenStudy (anonymous):

and the y value you should be able to see from looking at it.

OpenStudy (anonymous):

so it would be (3,4) x=-3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

just (3,4) not x=-3

OpenStudy (anonymous):

vertex/turning point is at (3,4)

OpenStudy (anonymous):

o sorry i need axis of symmetry too

OpenStudy (anonymous):

oh k. For the y-intercept, make x=0 in the equation for the x-intercept, make y=0 in the equation then solve the equation.

OpenStudy (anonymous):

o 3

OpenStudy (anonymous):

sorry, I do not understand what you just typed.

OpenStudy (anonymous):

i meant x=3 lol sorry

OpenStudy (anonymous):

let me solve it now, I'm pretty sure that is incorrect though, there should be two answers, give me a second.

OpenStudy (anonymous):

so you want to find the x-intercept I believe and to do so we have to make y=0 y=3(x-3)^2 + 4 0=3(x-3)^2 + 4 3(x^2-6x+9) + 4 = 0 3x^2 -18x + 27 + 4 = 0 3x^2 -18x + 31 = 0 here you can use the quadratic formula where a=3, b=-18 and c=31 \[ x = \frac{-b \pm \sqrt{b2-4ac}}{2a}\]

OpenStudy (anonymous):

and if you do that, you actually get no answer because there is in fact no x-intercept.

OpenStudy (anonymous):

you can try plot the graph on your calculator and you will see that the turning point as found before is at (3,4) and so that is the lowest point and it doesn't go past there, so clearly there is no x-intercept.

OpenStudy (anonymous):

got it thanks

OpenStudy (anonymous):

no problem, medal thanks :)